In 2008, Starbucks announced that they would be closing 600 US stores. Harvard Business Review points out that in this situation in order to keep up, Starbucks should improve their service. Starbuck founder Howard Schultz, than introduce a new concept to reduce customer waiting time. He believed that the customer will wait no longer than 400 seconds. Based on sample of 40 customers, the average amount of waiting time is 370 seconds and the sample standard deviation is 79. If the confidence level is 95%,

  1. Formulate hypotheses that can be used to challenge the validity of the star buck's founder claim
  2. Calculate the test statistics
  3. Find the p- value Based on the sample,

d. Test the hypotheses! Can we say the claim is correct? What is your explanation?
Answer #1

(a)
H0: Null Hypothesis: μ> 400

HA:Alternative Hypothesis:mu < 400

(b)

SE = s/sqrt{n}

= 79/V40 = 12.4910

Test statistic is:

t = (370 - 400)/12.4910= - 2.4017

So,

Test statistic = - 2.4017

(c)

t score = - 2.4017

ndf = 40 - 1 = 39

One Tail - Left Side Test

By Technology, P - Value= 0.0106

So,

P - Value = 0.0105

(d) Since P - value = 0.0106 is less than alpha = 0.05, the differenc is significant. Rejectnull hypothesis.

Conclusion:
The data support the claim that customer will wait no longer that 400 seconds.